### v1

@@ 1 Introduction
The comparison of heterogeneous data distributions is a fundamental task in computer vision, computational biology and machine learning. Most existing approaches rely on using a suitable ground cost function such as an available metric. A classic example is the Wasserstein distance which seeks an optimal transport (OT) between two given distributions. However, often the given distributions are in heterogeneous spaces, where a readily available ground cost function between these spaces does not generally exist. Additional effort may be required to learn appropriate cost functions Cuturi & Avis (2014); Heitz et al. (2021). However, even if there exists a natural embedding into a canonical joint metric space, this metric may not be suited to accurately gauge differences in their samples. Examples of such heterogeneous settings are, e.g., the comparison of graph- or mesh-valued data such as 3d shapes or manifolds. This paper introduces a novel framework based on OT which enables the joint comparison and visualization of heterogeneous datasets by optimally transferring them into an a-priori fixed metric space, see Figure 3 for an illustrative example.
@@ 1.1 Previous Work
Optimal Transport with Invariances. Optimal transport and Wasserstein distances enable comparisons of point clouds in metric spaces. However, classic OT lacks invariance to important transformations like rotations in Euclidean space. To address this, Gromov–Wasserstein (GW) distances are introduced which allow for the comparison of measures from distinct metric spaces Mémoli (2011); Sturm (2006). In Alaya et al. (2022), an approximation of GW distances is obtained by jointly embedding measures into Euclidean spaces. Furthermore, projection- and subspace-robust Wasserstein-2 distances are introduced in Paty & Cuturi (2019). Another approach incorporates invariance to Euclidean isometries via the Wasserstein Procrustes problem Grave et al. (2019). This is extended to account for linear operators with bounded Schatten norms in Alvarez-Melis et al. (2019) and to Gaussian mixture applications in Salmona et al. (2024). As outlined in Section 5, our paper extends such invariant OT to non-Euclidean domains.
Joint Dimensionality Reduction. Dimensionality reduction is a core topic in machine learning enabling visualization and clustering by finding optimal low-dimensional representations of data. Classical methods like principal component analysis (PCA) Greenacre et al. (2022) and multidimensional scaling (MDS) Carroll & Arabie (1998) preserve large variations or pairwise distances, but fail on nonlinear manifolds Alaya et al. (2022); Deng et al. (2024). Nonlinear approaches, e.g., locally linear embedding Roweis & Saul (2000), probabilistic models, e.g., t-distributed stochastic neighbor embedding (t-SNE) Van der Maaten & Hinton (2008), and deep learning methods, e.g., variational autoencoders (VAEs), Kingma et al. (2019) address these challenges. As an extension, several recent methods focus on joint embeddings of heterogeneous data. Here, we are given data on two incompatible domains and are interested in simultaneous embedding. The manifold-aligning generative adversarial network Amodio & Krishnaswamy (2018) employs a generative adversarial network for domain alignment. Maximum mean discrepancy (MMD) manifold-alignment Liu et al. (2019) balances an MMD and a distortion term. UnionCom Cao et al. (2020) leverages the generalized unsupervised manifold alignment (GUMA) Cui et al. (2014). Single-cell alignment with optimal transport (SCOT) Demetci et al. (2022) and the partial manifold alignment algorithm Cao et al. (2022) employ GW distances. Finally, the recently proposed joint multidimensional scaling (JMDS) Chen et al. (2023) algorithm combines MDS with OT for the joint embedding of two datasets into a shared Euclidean space. Complementing this and the work on non-Euclidean embeddings in McInnes et al. (2018); Deng et al. (2024), we propose a joint embedding into arbitrary metric spaces based on GW distances.
@@ 1.2 Contribution
2. We prove that our functional has a minimizer. Further, if its regularization parameter goes to infinity, our functional approaches the so-called “embedded Wasserstein distance” Salmona et al. (2024). In this sense, we also refer to our model as the “relaxed embedded Wasserstein distance”.
@@ 2 OT-Based Distances
Given a compact metric space $(Z,d_{Z})$, we denote by $\mathcal{P}(Z)$ the space of probability measures defined on the Borel-$\sigma$-algebra induced by the distance $d_{Z}$. For two compact metric spaces $Z_{1}$ and $Z_{2}$, the push-forward measure of $\mu\in\mathcal{P}(Z_{1})$ under a measurable map $T\colon Z_{1}\to Z_{2}$ is denoted by $T_{\sharp}\mu\coloneqq\mu\circ T^{-1}\in\mathcal{P}(Z_{2})$. The set of transport plans between two measures $\mu_{1}\in\mathcal{P}(Z_{1})$ and $\mu_{2}\in\mathcal{P}(Z_{2})$ is given by
@@
Sturm’s GW distance Sturm (2006) between two mm-spaces $\mathbb{X}_{i}$, $i=1,2$, is defined by
@@
Mémoli proposed a different GW distance, which is numerically more appealing than Sturm’s construction. For two mm-spaces $\mathbb{X}_{i}$, $i=1,2$, Mémoli’s GW distance Mémoli (2011) is defined by
@@
The problem seeks a transport between $\mathbb{X}_{1}$ and $\mathbb{X}_{2}$ which minimizes the overall pairwise distance distortion, see Figure 4(b). Both distances—SGW and GW— define a metric on the equivalence classes of mm-spaces. Furthermore, both distances induce the same topology, but SGW turns these equivalence classes into a complete metric space, while GW does not Sturm (2006); Mémoli (2011). The later can be alleviated by embedding into a larger space and extending the GW definition accordingly Sturm (2023).
@@ 3 Unbalanced OT With GW Penalization
On a certain subspace of equivalence classes of mm-spaces, this defines a metric, which we call embedded Wasserstein metric. For the specific case of Euclidean spaces $X_{1}$, $X_{2}$, and $Z$, this reduces to the Wasserstein Procrustes problem Grave et al. (2019) and to the “embedded Wasserstein metric” in Salmona et al. (2024). We adopt this name.
@@
While EW relies on appropriate isometries, the following relaxation enables to handle arbitrary mm-spaces $\mathbb{X}_{i}$, $i=1,2$. For $\lambda>0$, we consider the following GW penalized unbalanced OT problem
@@
If $\alpha$ is a solution of (14), then its projections (13) are solutions of (10) and in particular $\pi$ is a solution of (8). Conversely, any solution of (8) can be expressed in this way.
@@
Figure 5: Illustration of our multi-marginal transport problem.
@@ 4 Bi-Convex Relaxation
At its core, the computation of EWλ in (14) requires the solution of a quadratic optimization problem. Similar formulations appear, for instance, in the computation of the GW distance Peyré et al. (2016); Séjourné et al. (2021), in the multi-marginal GW setting Beier et al. (2023), and in CO-OT Vayer et al. (2020b). All of these quadratic OT problem have in common that they can be numerically solved using a block-coordinate descent on their bi-convex relaxations. In the following, we adapt this approach to our multi-marginal, unbalanced transport problem (14).
@@
The cost function additively decouples into three partial cost functions solely depending on two coordinates of $c_{1}$. Unbalanced, multi-marginal OT formulations of this kind are treated in Beier et al. (2022), where, relying on an entropic regularization, a multi-marginal Sinkhorn scheme is proposed. Mathematically, this means that, for a regularization parameter $\varepsilon>0$, we approximate the minimizer of (22) by
@@
4:     Update $\alpha_{2}$ using the multi-marginal Sinkhorn scheme in Beier et al. (2022) with $c_{1}$
@@
6:     Update $\alpha_{1}$ using the multi-marginal Sinkhorn scheme in Beier et al. (2022) with $c_{2}$
@@ 5 Discrete EWλ and JMDS
The distance EW in (6) generalizes Wasserstein Procrustes Grave et al. (2019) to non-Euclidean spaces, while EWλ in (8) is closely related to the JMDS model Chen et al. (2023). In this section, we explain the relation.
@@
where $i=1,2$. Since the supports of the measures are fixed, we can restrict ourselves to the weight matrices $\boldsymbol{\mu}_{i}\in\Delta_{m}$, $\boldsymbol{\xi}_{i}\in\Delta_{n_{i}}$, $\boldsymbol{\pi}\in\Pi(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2})$ and $\boldsymbol{\gamma}_{i}\in\Pi(\boldsymbol{\xi}_{i},\boldsymbol{\mu}_{i})$, $i=1,2$ in the corresponding probability simplicies $\Delta$. Then EWλ in (12) becomes the discrete minimization problem
$\operatorname{EW}_{\lambda}(X_{1},X_{2})=\min_{\begin{subarray}{c}\boldsymbol{\mu_{i}}\in\Delta_{m}\\ i=1,2\end{subarray}}\min_{\boldsymbol{\pi}\in\Pi(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2})}\min_{\begin{subarray}{c}\boldsymbol{\gamma}_{i}\in\Pi(\boldsymbol{\xi}_{i},\boldsymbol{\mu}_{i})\\ i=1,2\end{subarray}}\;$ (27)
$\big{\langle}\boldsymbol{\pi},\boldsymbol{D}^{2}_{Z,Z}\big{\rangle}+\lambda\big{(}G_{X_{1},Z}(\boldsymbol{\gamma}_{1})+G_{X_{2},Z}(\boldsymbol{\gamma}_{2})\big{)},$ (28)
where $\boldsymbol{D}^{2}_{Z,Z}\coloneqq\big{(}d_{Z}^{2}(z^{j},z^{k})\big{)}_{j,k=1}^{m}$ and, for $i=1,2$,
$G_{X_{i},Z}(\boldsymbol{\gamma}_{i})$ (29)
$\coloneqq\sum_{j,k=1}^{n_{i},m}\sum_{r,s=1}^{n_{i},m}\gamma_{i}^{s,t}\gamma_{i}^{j,k}\big{(}d_{X_{1}}(x_{i}^{j},x_{i}^{r})-d_{Z}(z^{k},z^{s})\big{)}^{2}.$ (30)
@@
$\operatorname{JMDS}_{\lambda}(X_{1},X_{2})\coloneqq\min_{\mathcal{Z}_{1},\mathcal{Z}_{2}}\min_{\boldsymbol{\pi}\in\Pi(\frac{1}{n_{1}}\boldsymbol{1}_{n_{1}},\frac{1}{n_{2}}\boldsymbol{1}_{n_{2}})}\;$ (32)
$\big{\langle}\pi,\boldsymbol{D}^{2}_{\mathcal{Z}_{1},\mathcal{Z}_{2}}\big{\rangle}+\lambda\big{(}G_{1}(\mathcal{Z}_{1})+G_{2}(\mathcal{Z}_{2})\big{)},$ (33)
where $\boldsymbol{D}^{2}_{\mathcal{Z}_{1},\mathcal{Z}_{2}}\coloneqq\big{(}d^{2}(z_{1}^{j},z_{2}^{k})\big{)}_{j,k=1}^{n_{1},n_{2}}$ and, for $i=1,2$,
@@
This is exactly the functional proposed as JMDS in Chen et al. (2023). Note that the Wasserstein Procrustes model is given by
$\min_{\boldsymbol{\pi}\in\Pi(\frac{1}{n_{1}}\boldsymbol{1}_{n_{1}},\frac{1}{n_{2}}\boldsymbol{1}_{n_{2}})}\min_{Q\in\operatorname{SO}(d)}\big{\langle}\pi,\boldsymbol{D}^{2}_{Q\mathcal{Z}_{1},\mathcal{Z}_{2}}\big{\rangle}.$ (36)
@@
To summarize: 1. Our discrete EWλ fixes the support of the marginals $\mu_{i}$ in (12) which results in the optimization over the weights, where JMSD fixes the weights of the $\mu_{i}$ and optimizes over the supports. 2. Fixing the support has the advantage that we can work with arbitrary metric spaces $(Z,d_{Z})$, while the “free support” approach of JMDS is restricted to the Euclidean space. 3. The minimization problems have to be tackled by completely different optimization algorithms, namely an unbalanced multimarginal Sinkhorn algorithm for the block-coordinate descent in Algorithm 1 for EWλ and the so-called SMACOF method combined with Wasserstein Procrustes minimizations Chen et al. (2023) for JMDS.

### v2

@@ 1 Introduction
The comparison of heterogeneous data distributions is a fundamental task in computer vision, computational biology and machine learning. Most existing approaches rely on using a suitable ground cost function such as an available metric. A classic example is the Wasserstein distance which seeks an optimal transport (OT) between two given distributions. However, often the given distributions are in heterogeneous spaces, where a readily available ground cost function between these spaces does not generally exist. Additional effort may be required to learn appropriate cost functions (Cuturi & Avis, 2014; Heitz et al., 2021). However, even if there exists a natural embedding into a canonical joint metric space, this metric may not be suited to accurately gauge differences in their samples. Examples of such heterogeneous settings are, e.g., the comparison of graph- or mesh-valued data such as 3d shapes or manifolds. This paper introduces a novel framework based on OT which enables the joint comparison and visualization of heterogeneous datasets by optimally transferring them into an a-priori fixed metric space, see Figure 3 for an illustrative example.
@@ 1.1 Previous Work
Optimal Transport with Invariances
Classical Wasserstein distances allow to compare measures on a common metric space. In specific tasks in geometry processing like shape matching, the considered measures, however, live on distinct metric spaces that incorporate the different geodesic distances on the given shapes. To address this, Gromov–Wasserstein (GW) distances are introduced (Mémoli, 2011; Sturm, 2006). In (Alaya et al., 2022), an approximation of GW distances is obtained by jointly embedding measures into Euclidean spaces. Furthermore, projection- and subspace-robust Wasserstein-2 distances are introduced in (Paty & Cuturi, 2019). Another approach incorporates invariance to Euclidean isometries via the Wasserstein Procrustes problem (Grave et al., 2019). This is extended to account for linear operators with bounded Schatten norms in (Alvarez-Melis et al., 2019) and to Gaussian mixture applications in (Salmona et al., 2024). As outlined in Section 5, our paper extends such invariant OT to non-Euclidean domains.
Joint Dimensionality Reduction
Dimensionality reduction is a core topic in machine learning enabling visualization and clustering by finding optimal low-dimensional representations of data. Classical methods like principal component analysis (PCA) (Greenacre et al., 2022) and multidimensional scaling (MDS) (Carroll & Arabie, 1998) preserve large variations or pairwise distances, but fail on nonlinear manifolds (Alaya et al., 2022; Deng et al., 2024). Nonlinear approaches, e.g., locally linear embedding (Roweis & Saul, 2000), probabilistic models, e.g., t-distributed stochastic neighbor embedding (t-SNE) (Van der Maaten & Hinton, 2008), and deep learning methods, e.g., variational autoencoders (VAEs), (Kingma et al., 2019) address these challenges. As an extension, several recent methods focus on joint embeddings of heterogeneous data. Here, we are given data on two incompatible domains and are interested in simultaneous embedding. The manifold-aligning generative adversarial network (Amodio & Krishnaswamy, 2018) employs a generative adversarial network for domain alignment. Maximum mean discrepancy (MMD) manifold-alignment (Liu et al., 2019) balances an MMD and a distortion term. UnionCom (Cao et al., 2020) leverages the generalized unsupervised manifold alignment (GUMA) (Cui et al., 2014). Single-cell alignment with optimal transport (SCOT) (Demetci et al., 2022) and the partial manifold alignment algorithm (Cao et al., 2022) employ GW distances. Finally, the recently proposed joint multidimensional scaling (JMDS) (Chen et al., 2023) algorithm combines MDS with OT for the joint embedding of two datasets into a shared Euclidean space. Complementing this and research on non-Euclidean embeddings (McInnes et al., 2018; Deng et al., 2024) and connections between MDS and GW (Van Assel et al., 2024; Clark et al., 2025), we propose a joint embedding into arbitrary metric spaces based on factored transport plans (Forrow et al., 2019) and GW distances.
@@ 1.2 Contribution
2. We prove that our functional has a minimizer. Further, if its regularization parameter goes to infinity, our functional approaches the so-called “embedded Wasserstein distance” (Salmona et al., 2024). In this sense, we also refer to our model as the “relaxed embedded Wasserstein distance”.
@@ 2 OT-Based Distances
Given a compact metric space $(Z,d_{Z})$, we denote by $\mathcal{P}(Z)$ the space of probability measures defined on the Borel-$\sigma$-algebra induced by the distance $d_{Z}$. For two compact metric spaces $Z_{1}$ and $Z_{2}$, the push-forward measure of $\mu\in\mathcal{P}(Z_{1})$ under a measurable map $T\colon Z_{1}\to Z_{2}$ is denoted by
$T_{\sharp}\mu\coloneqq\mu\circ T^{-1}\in\mathcal{P}(Z_{2}).$ (1)
The set of transport plans between two measures $\mu_{1}\in\mathcal{P}(Z_{1})$ and $\mu_{2}\in\mathcal{P}(Z_{2})$ is given by
@@
Sturm’s GW distance (Sturm, 2006) between two mm-spaces $\mathbb{X}_{i}$, $i=1,2$, is defined by
@@
Mémoli proposed a different GW distance, which is numerically more appealing than Sturm’s construction. For two mm-spaces $\mathbb{X}_{i}$, $i=1,2$, Mémoli’s GW distance (Mémoli, 2011) is defined by
@@
The problem seeks a transport between $\mathbb{X}_{1}$ and $\mathbb{X}_{2}$ which minimizes the overall pairwise distance distortion, see Figure 4(b). Both distances—SGW and GW— define a metric on the equivalence classes of mm-spaces. Furthermore, both distances induce the same topology, but SGW turns these equivalence classes into a complete metric space, while GW does not (Sturm, 2006; Mémoli, 2011). The later can be alleviated by embedding into a larger space and extending the GW definition accordingly (Sturm, 2023).
@@ 3 Unbalanced OT With GW Penalization
On a certain subspace of equivalence classes of mm-spaces, this defines a metric, which we call embedded Wasserstein metric. For the specific case of Euclidean spaces $X_{1}$, $X_{2}$, and $Z$, this reduces to the Wasserstein Procrustes problem (Grave et al., 2019) and to the “embedded Wasserstein metric” in (Salmona et al., 2024). We adopt this name.
@@
While EW relies on appropriate isometries, the following relaxation enables us to handle arbitrary mm-spaces $\mathbb{X}_{i}$, $i=1,2$. For $\lambda>0$, we consider the following GW penalized unbalanced OT problem
@@
If there exist no isometric embeddings $\operatorname*{supp}\xi_{i}\hookrightarrow Z$, the limit of $\operatorname{EW}_{\lambda}$ as $\lambda\to\infty$ yields best possible approximations of $\mathbb{X}_{i}$ on the given metric space $(Z,d_{Z})$ with respect to the GW distance. More precisely, a GW approximation of an mm-space $\mathbb{X}$ on the metric space $(Z,d_{Z})$ is a minimizer $\zeta\in\mathcal{P}(Z)$ of
$\operatorname{GWA}(\mathbb{X})\coloneqq\min_{\zeta\in\mathcal{P}(Z)}\operatorname{GW}^{2}(\mathbb{X},(Z,d_{Z},\zeta)).$ (10)
For discrete mm-spaces, GW approximations are studied in (Clark et al., 2025) and are closely related to MDS.
Proposition 3.5.
Let $\mathbb{X}_{i}$, $i=1,2$, be two mm-spaces and let $(Z,d_{Z})$ be a metric space. Let $(\lambda_{n})_{n\in\mathbb{N}}$ be a sequence with $\lambda_{n}\to\infty$ as $n\to\infty$. Then any sequence $(\pi_{n})_{n\in\mathbb{N}}$ of minimizers of $\operatorname{EW}_{\lambda_{n}}(\mathbb{X}_{1},\mathbb{X}_{2})$ converges weakly, up to a subsequence, to $\pi\in\Pi(\zeta_{1},\zeta_{2})$, where $\zeta_{i}\in\mathcal{P}(Z)$ is a GW approximation of $\mathbb{X}_{i}$.
On the other side, if $\lambda\to 0$, the influence of the Wasserstein term in $\operatorname{EW}_{\lambda}$ increases. Figuratively, the marginals $P_{1,\sharp}\pi$ and $P_{2,\sharp}\pi$ of the minimizers $\pi$ of $\operatorname{EW}_{\lambda}$ have to become more similar if $\lambda$ goes to zero. In the limit case, when both marginals coincide, we obtain a fixed-support GW barycenter (Beier et al., 2023). In detail, a fixed-support GW barycenter on the metric space $(Z,d_{Z})$ is a minimizer $\zeta\in\mathcal{P}(Z)$ of
$\operatorname{GWB}(\mathbb{X}_{1},\mathbb{X}_{2})\coloneqq\min_{\zeta\in\mathcal{P}(Z)}\sum_{i=1}^{2}\operatorname{GW}^{2}(\mathbb{X}_{i},(Z,d_{Z},\zeta)).$ (11)
Proposition 3.6.
Let $\mathbb{X}_{i}$, $i=1,2$, be two mm-spaces and let $(Z,d_{Z})$ be a metric space. Let $(\lambda_{n})_{n\in\mathbb{N}}$ be a sequence with $\lambda_{n}\to 0$ as $n\to\infty$. Then any sequence $(\pi_{n})_{n\in\mathbb{N}}$ of minimizers of $\operatorname{EW}_{\lambda_{n}}(\mathbb{X}_{1},\mathbb{X}_{2})$ converges weakly, up to a subsequence, to $(\operatorname{Id},\operatorname{Id})_{\sharp}\zeta$, where $\zeta\in\mathcal{P}(Z)$ is a fixed-support GW barycenter.
@@
Figure 5: Illustration of our multi-marginal transport problem. The colors are in line with Figure 4: quantities with respect to $\mathbb{X}_{1}$ are red, and quantities related to $\mathbb{X}_{2}$ are blue.
@@
If $\alpha$ solves (17), then its projections (16) are solutions of (13) and in particular $\pi$ is a solution of (9). Conversely, any solution of (9) can be expressed this way.
@@ 4 Bi-Convex Relaxation
At its core, the computation of EWλ in (17) requires the solution of a quadratic optimization problem. Similar formulations appear, for instance, in the computation of the GW distance (Peyré et al., 2016; Séjourné et al., 2021), in the multi-marginal GW setting (Beier et al., 2023), and in CO-OT (Vayer et al., 2020b). All of these quadratic OT problem have in common that they can be numerically solved using a block-coordinate descent on their bi-convex relaxations. In the following, we adapt this approach to our multi-marginal, unbalanced transport problem (17).
@@
The cost function additively decouples into three partial cost functions solely depending on two coordinates of $c_{1}$. Unbalanced, multi-marginal OT formulations of this kind are treated in (Beier et al., 2022), where, relying on an entropic regularization, a multi-marginal Sinkhorn scheme is proposed. Mathematically, this means that, for a regularization parameter $\varepsilon>0$, we approximate the minimizer of (25) by
@@
4:     Update $\alpha_{2}$ using the multi-marginal Sinkhorn scheme in (Beier et al., 2022) with $c_{1}$
@@
6:     Update $\alpha_{1}$ using the multi-marginal Sinkhorn scheme in (Beier et al., 2022) with $c_{2}$
@@ 5 Discrete EWλ and JMDS
The distance EW in (7) generalizes Wasserstein Procrustes (Grave et al., 2019) to non-Euclidean spaces, while EWλ in (9) is closely related to the JMDS model (Chen et al., 2023). In this section, we explain the relation.
@@
where $i=1,2$. Since the supports of the measures are fixed, we can restrict ourselves to the weight matrices $\bm{\mu}_{i}\in\Delta_{m}$, $\bm{\xi}_{i}\in\Delta_{n_{i}}$, $\bm{\pi}\in\Pi(\bm{\mu}_{1},\bm{\mu}_{2})$ and $\bm{\gamma}_{i}\in\Pi(\bm{\xi}_{i},\bm{\mu}_{i})$, $i=1,2$ in the corresponding probability simplicies $\Delta$. Then EWλ in (15) becomes the discrete minimization problem
$\operatorname{EW}_{\lambda}(X_{1},X_{2})=\min_{\begin{subarray}{c}\bm{\mu_{i}}\in\Delta_{m}\\ i=1,2\end{subarray}}\min_{\bm{\pi}\in\Pi(\bm{\mu}_{1},\bm{\mu}_{2})}\min_{\begin{subarray}{c}\bm{\gamma}_{i}\in\Pi(\bm{\xi}_{i},\bm{\mu}_{i})\\ i=1,2\end{subarray}}\;$ (30)
$\big{\langle}\bm{\pi},\bm{D}^{2}_{Z,Z}\big{\rangle}+\lambda\big{(}G_{X_{1},Z}(\bm{\gamma}_{1})+G_{X_{2},Z}(\bm{\gamma}_{2})\big{)},$ (31)
where $\bm{D}^{2}_{Z,Z}\coloneqq\big{(}d_{Z}^{2}(z^{j},z^{k})\big{)}_{j,k=1}^{m}$ and, for $i=1,2$,
$G_{X_{i},Z}(\bm{\gamma}_{i})$ (32)
$\coloneqq\sum_{j,k=1}^{n_{i},m}\sum_{r,s=1}^{n_{i},m}\gamma_{i}^{s,k}\gamma_{i}^{j,k}\big{(}d_{X_{1}}(x_{i}^{j},x_{i}^{r})-d_{Z}(z^{k},z^{s})\big{)}^{2}.$ (33)
@@
$\operatorname{JMDS}_{\lambda}(X_{1},X_{2})\coloneqq\min_{\mathcal{Z}_{1},\mathcal{Z}_{2}}\min_{\bm{\pi}\in\Pi(\frac{1}{n_{1}}\bm{1}_{n_{1}},\frac{1}{n_{2}}\bm{1}_{n_{2}})}\;$ (35)
$\big{\langle}\pi,\bm{D}^{2}_{\mathcal{Z}_{1},\mathcal{Z}_{2}}\big{\rangle}+\lambda\big{(}G_{1}(\mathcal{Z}_{1})+G_{2}(\mathcal{Z}_{2})\big{)},$ (36)
where $\bm{D}^{2}_{\mathcal{Z}_{1},\mathcal{Z}_{2}}\coloneqq\big{(}d^{2}(z_{1}^{j},z_{2}^{k})\big{)}_{j,k=1}^{n_{1},n_{2}}$ and, for $i=1,2$,
@@
This is exactly the functional proposed as JMDS in (Chen et al., 2023). Note that the Wasserstein Procrustes model is given by
$\min_{\bm{\pi}\in\Pi(\frac{1}{n_{1}}\bm{1}_{n_{1}},\frac{1}{n_{2}}\bm{1}_{n_{2}})}\min_{Q\in\operatorname{SO}(d)}\big{\langle}\pi,\bm{D}^{2}_{Q\mathcal{Z}_{1},\mathcal{Z}_{2}}\big{\rangle}.$ (39)
@@
To summarize: 1. Our discrete EWλ fixes the support of the marginals $\mu_{i}$ in (15) which results in the optimization over the weights, where JMSD fixes the weights of the $\mu_{i}$ and optimizes over the supports. 2. Fixing the support has the advantage that we can work with arbitrary metric spaces $(Z,d_{Z})$, while the “free support” approach of JMDS is restricted to the Euclidean space. 3. The minimization problems have to be tackled by completely different optimization algorithms, namely an unbalanced multimarginal Sinkhorn algorithm for the block-coordinate descent in Algorithm 1 for EWλ and the so-called SMACOF method combined with Wasserstein Procrustes minimizations (Chen et al., 2023) for JMDS.